OpenStudy (anonymous):

Solve: 6x − 2(x + 1) = 3x − 10 8 −8 −12 12

OpenStudy (anonymous):

x=-8

OpenStudy (anonymous):

-12 :)

OpenStudy (anonymous):

im confused which one is it

OpenStudy (anonymous):

How about you plug them in and see if they work? .. it's rather easy to check your answer here.

OpenStudy (anonymous):

\[6x-2(x+1)=3x-10 is equal \to 6x-2x-2=3x-10 then 4x-2=3x-10 then 4x-3x=-10+2 then x=-8\]

OpenStudy (anonymous):

you do -2*x which is -2x and then -2*1 which is -2 then combine 6x and -2x and get 4x next subtract 3x from 4x and you get 1x then -2 from 10 and get -12 then divide -12 by 1x and you get -12

OpenStudy (anonymous):

10 - -2 isn't -12... it's -8.

OpenStudy (anonymous):

-10 **

OpenStudy (anonymous):

ohh i see what i did i put minus instead of negative , thanks :) @mustry was correct

OpenStudy (anonymous):

thankx much am in Africa